If a rectangle has a perimeter of 206 ft, and the length is 2 less than 6 times it's width, what is the length and width?
Answer by CubeyThePenguin(3113)

width = w
length = 6w - 2

2(w) + 2(6w - 2) = 206
2w + 12w - 4 = 206
14w = 210

width = w = 15
length = 6w - 2 = 88